# A non-degeneracy theorem for interacting electrons in one dimension

### v1

@@ 1. Introduction
acting on the space of antisymmetric wave-functions $\wedge^{N}\mathrm{L}^{2}(I)$, where $I=(0,1)$ (or $\mathbb{R}$), and $v$ and $w$ belongs to suitable class of distributions. Such operators are used to describe interacting electrons living in one-dimensional space and are therefore ubiquitous in quantum mechanics.
@@
In fact, our main motivation for studying the interacting operator $H_{N}(v,w)$ stems from recent advances [SPR+24, Cor25] towards establishing a mathematically rigorous formulation of Kohn-Sham Density Functional Theory (DFT), a highly successful and widely used approach to study the electronic structure of materials. More precisely, the class of distributional potentials studied here appeared in recent developments [SPR+24] concerning the $v$-representability problem, which, roughly speaking, consists in characterizing the set of all possible single-particle ground state densities of operators of the form $H_{N}(v,w)$ for a given class of potentials $v$. While a detailed discussion on the foundation of DFT and the $v$-representability problem goes beyond the scope of the current paper, we emphasize that the results obtained here will be used to provide a completely rigorous formulation of Kohn-Sham DFT for electrons in one-dimension in a follow-up article.
@@ 2.1. Notation
We denote by $\mathrm{H}^{1}(I)$ the Sobolev space of functions $f\in\mathrm{L}^{2}(I)$ with weak derivative $\partial_{x}f\in\mathrm{L}^{2}(I)$. More generally, for $1\leq p\leq\infty$ and open $\Omega\subset\mathbb{R}^{d}$, we denote by $\mathrm{W}^{1,p}(\Omega)$ the usual Sobolev spaces of functions in $\mathrm{L}^{p}(\Omega)$ with weak gradient in $\mathrm{L}^{p}(\Omega)$, and by $\mathrm{W}^{-1,q}(\Omega)$, where $11/q+1/p=$ 1, the dual space of $\mathrm{W}^{1,p}(\Omega)$. In addition, we denote by $\mathrm{H}^{1/2}(\partial\Omega)$ the standard $1/2$-Sobolev (or Besov) space along the boundary $\partial\Omega$.
For a given closed subspace $L\subset\mathrm{H}^{1/2}(\partial\Omega)$, we denote by $H^{1}_{L}(\Omega)$ the Sobolev spaces of functions with trace on $L$, i.e.,
$\mathrm{H}^{1}_{L}(\Omega)\coloneqq\{f\in\mathrm{H}^{1}(\Omega):\gamma f\in L\},$
where $\gamma:\mathrm{H}^{1}(\Omega)\rightarrow\mathrm{H}^{1/2}(\partial\Omega)$ is the Dirichlet (boundary) trace operator.
@@
$\mathcal{V}\coloneqq\{v\in\mathrm{H}^{-1}(I):v(\varphi)\in\mathbb{R}\quad\mbox{for any real-valued $\varphi\in\mathrm{H}^{1}(I)$}\}.$
@@
$\mathcal{W}\coloneqq\{w\in\mathrm{W}^{-1,q}(I_{2}):\mbox{for some}q>2\quad\mbox{and}\quad w(\varphi)\in\mathbb{R}\quad\mbox{for any real-valued $\varphi\in\mathrm{W}^{1,p}(I_{2})$}\}.$
@@
For $v\in\mathcal{V}$, $w\in\mathcal{W}$, and $L\subset\mathrm{H}^{1/2}(\partial I_{N})$ as above, we denote by $H_{N}^{L}(v,w)$ the $N$-particle Hamiltonian
$H_{N}^{L}(v,w)=-\Delta+\sum_{i\neq j}^{N}w(x_{i},x_{j})+\sum_{j=1}^{N}v(x_{i})\quad\mbox{acting on $\mathcal{H}_{N}$ with boundary conditions in $L$.}$
More precisely, $H_{N}^{L}(v,w)$ is the unique self-adjoint operator associated to the sesquilinear form
$a_{v,w}(\Psi,\Phi)\coloneqq\int_{I_{N}}\overline{\nabla\Psi(x_{1},...,x_{N})}\bm{\cdot}\nabla\Phi(x_{1},...,x_{N})\mathrm{d}x_{1}...\mathrm{d}x_{N}+v\left(\rho_{\Psi,\Phi}\right)+w(\rho_{2,\Psi,\Phi})$ (2.1)
@@
$Q_{N}^{L}\coloneqq\mathcal{H}_{N}\cap\mathrm{H}^{1}_{L}(I_{N}),$
@@
and $\rho_{2,\Psi,\Phi}$ is the overlapping pair density
$\rho_{2,\Psi,\Phi}(x,y)\coloneqq N(N-1)\int_{I_{N-2}}\overline{\Psi(x,y,x_{3},...,x_{N})}\Phi(x,y,x_{3},...,x_{N})\mathrm{d}x_{3}...\mathrm{d}x_{N}.$ (2.3)
@@ 2.2. Main results
As a first result, we show that the ground-state of self-adjoint realizations of $H_{N}(v,w)$ with local boundary conditions is non-degenerate. Precisely, let $\Gamma\subset\partial\Omega$ and define
@@
Let $v\in\mathcal{V}$, $w\in\mathcal{W}$, $N\in\mathbb{N}$ and $\Gamma\subset\partial\Omega$. Then the operator $H_{N}(v,w;\Gamma)$ defined as the self-adjoint realization of (2.1) with form domain $\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}$ has a unique ground-state $\Psi$ and $\Psi\neq 0$ almost everywhere.
@@
$\mathrm{H}^{1}_{\alpha}(I_{N})\coloneqq\left\{\Psi\in\mathrm{H}^{1}(I_{N}):(\gamma\Psi)\rvert_{I_{k-1}\times\{0\}\times I_{N-k}}-\alpha(\gamma\Psi)\rvert_{I_{k-1}\times\{1\}\times I_{N-k}}=0\quad\mbox{for any $1\leq k\leq N$}\right\}.$
@@
Then the self-adjoint realization $H_{N}^{\alpha}(v,w)$ of (2.1) with form domain $\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}$ has a unique ground state $\Psi$ and $\Psi\neq 0$ almost everywhere.
@@
Theorem 2.3 shows that periodic systems are non-degenerate for an odd number of particles, while anti-periodic systems are non-degenerate for an even number of particles. This result is optimal in the sense that there are explicit examples where the ground-state is non-degenerate if and only if condition (2.4) holds. For instance, the simplest example is the free periodic and anti-periodic Laplacians, where the ground-state is non-degenerate if and only if $N$ is odd respectively even. Moreover, one can also construct examples of periodic systems where the ground-state is non-degenerate for any given number of particles $N$; thus an improvement of the form ”the system is at least two-fold degenerate for even/odd number of particles” is also not possible.
@@
    (Many-body interactions) We can allow a larger class of distributional interaction potentials $w$. In fact, Theorem 2.1 holds for $M$-body interactions $w_{M}$ with arbitrary $M\in\mathbb{N}$, provided that $w_{M}$ lies in suitable dual Sobolev spaces (see Section 4.3). Notably, the case of two-body 3D Coulomb interactions, which could be seen as a ”six-body” interaction of the form
@@
    (More general boundary conditions) We can further extend Theorem 2.1 to a larger class of self-adjoint realizations of $H_{N}(v,w)$. Roughly speaking, the maximal class of spaces for which our proof works is given by all $\mathrm{H}^{1}_{L}(I_{N})$ where the restricted space
    $\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{L}(I_{N})\cap\mathcal{H}_{N}\}\quad\mbox{where}\quad S_{N}=\{(x_{1},...,x_{N})\in I_{N}:x_{1}<x_{2}...<x_{N}\}$
@@
Interestingly, Theorems 2.1 and 2.3 can be applied to obtain information on the whole spectrum of single-particle operators. More precisely, let $v\in\mathcal{V}$, $L\subset\mathbb{C}^{2}$ be a (closed) subspace, and denote by $h_{L}(v)=-\Delta_{L}+v$ the self-adjoint operator associated to the form
$a_{v}(\psi,\psi)=\int_{I}|\nabla\psi(x)|^{2}\mathrm{d}x+v(|\psi|^{2})\quad\mbox{with form domain $\mathrm{H}^{1}_{L}(I)=\{\psi\in\mathrm{H}^{1}(I):(\psi(0),\psi(1))\in L\}$.}$ (2.6)
Then by applying Theorems 2.1 and 2.3 to self-adjoint realizations of the non-interacting operator $H_{N}(v,0)$, we obtain the following result.
@@
Let $v\in\mathcal{V}$ and $L=\{\beta=(\beta_{0},\beta_{1})\in\mathbb{C}^{2}:\beta_{0}-\alpha\beta_{1}=0\}$ for some $\alpha\in\mathbb{R}\setminus\{0\}$. Then $h_{L}(v)$ has discrete spectrum, all eigenfunctions are almost everywhere non-vanishing, and the eigenvalues $\{\lambda_{k}\}_{k\in\mathbb{N}}$ ordered in non-decreasing order and counting multiplicity satisfy
@@
Moreover, in the case where $L$ is one of the following sets
@@
Remark (Self-adjont realizations).
We remark that all local self-adjoint realizations of (2.6) on $C_{c}^{\infty}(I)$ are of the form $h_{L}(v)$ for some closed subspace $L\subset\mathbb{C}^{2}$ (see Propostion 3.9). Hence, Theorem 2.6 provides information on the spectrum of all self-adjoint extensions of $h_{L}(v)$ in $C_{c}^{\infty}(I)$ whose form domain is closed under complex conjugation.
@@
The result from Theorem 2.6 in the case of local boundary conditions (2.8) is, at least for regular potentials $v$, well-known in the theory of Sturm-Liouville operators, see, e.g., [Zet21, Theorem 1.9.2]. Nevertheless, the approach presented here seems to be completely different from previous proofs and has the important advantage that it works in the interacting case.
@@
Moreover, if $\alpha\in\mathbb{R}\setminus\{0\}$ satisfies (2.4) and $\lambda_{1}(\Gamma)$ and $\lambda_{1}(\Gamma^{\prime})$ denote respectively the lowest eigenvalues of the self-adjoint realizations with form domain $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}$ and $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma^{\prime})\cap\mathcal{H}_{N}$, then the same result holds.
@@ 2.3. Outline of the proofs
The proof of Theorems 2.1 and 2.3 is carried out in Section 5 and relies on two main steps. The first step consists in proving a Perron-Frobenius (PF) theorem for Schrödinger operators of the form $-\Delta+V$ for a large class of distributional potentials $V$ (see Theorem 5.3). In other words, we show that the ground-state of such operators is non-degenerate and strictly positive almost everywhere. This step relies on well-known semigroup techniques and a perturbative argument borrowed from Reed and Simon [RS78][Section XII.12] (see also [GJ70, FS75, Ges84] for similar arguments). Combining this result with a density result for dual Sobolev spaces (Lemma 3.11), we can then establish a PF theorem for Schrödinger operators of the form $H_{N}(v,w)=-\Delta+\sum_{j}v_{j}+\sum_{j\neq i}w_{i,j}$ under a suitable positivity assumption on their form domain. Unfortunately, these results only apply to the operator without Fermi statistics, i.e., without antisymmetry restrictions on the wave-function. Indeed, whenever antisymmetry is present the wave-function must have both positive and negative values and can not be almost everywhere strictly positive. In particular, as warned in [RS78, page 207], PF results cannot be applied to electronic systems.
@@
In Section 5.4 we present the proof of Theorem 2.6. This result is an immediate application of Theorems 2.1 and 2.3, and a technical lemma that connects the boundary conditions of the non-interacting operator $H_{N}(v,0)$ to the boundary conditions of the single-particle operator $h(v)=-\Delta+v$.
@@
The proof of Theorem 2.7 is inspired by ideas from previous works on similar eigenvalue inequalities [Fil05, GM09, BRS18, Roh21]. More precisely, the core idea of the proof is to apply the beautiful domain extension argument of Filonov [Fil05] to establish a unique continuation result along the boundary (cf. Theorem 6.1). Here, however, we face two fundamental problems when trying to apply this argument. First, it relies on unique continuation results which are not available in the case of distributional potentials. Second, and perhaps more critical, is the fact that the non-local boundary conditions are such that, roughly speaking, the faces of $I_{N}$ are ”glued” to each other, and therefore, no extension argument is possible.
@@
The first problem can be overcome by using the unique continuation property of the ground-state guaranteed by the PF theorem. Note that this property is only proven for the ground-state, which explains why we can only obtain inequalities for the lowest eigenvalue of $H_{N}(v,w)$. To deal with the second difficulty, we rely on a distributional definition of the Neumann trace (or normal derivative). While this definition is the same as in previous works, e.g., [GM09] where the Laplacian with non-local (Robin) boundary conditions is considered, we emphasize that the situation here is more delicate for two reasons. First, the non-local boundary conditions here are imposed directly in the form domain of the operator, as opposed to adding a non-local form along the boundary as done in [GM09]. Second, an eigenfunction $\Psi$ of $H_{N}(v,w)$ has, in general, no more regularity than $\mathrm{H}^{1}$. In particular, its Laplacian has no more regularity than $\mathrm{H}^{-1}$, and therefore, the usual definition of the weak Neumann trace, which requires $\Delta\Psi$ to be at least in $\mathrm{H}^{s}$ for $s>-1/2$, see [GM09, (3.17)], is not meaningful.
@@
Fortunately, we can still show that, despite this lack of regularity, a meaningful sense can be given to the Neumann trace of eigenfunctions along any open subset of the boundary where they vanish; this result crucially relies on the locality of the operator $H_{N}(v,w)$, a property that indirectly underlies many of the arguments used here. We then obtain an explicitly formula for this weak Neumann trace that could be new and of independent interest; see Lemma 6.4. By combining this formula with a positivity property of the boundary conditions, which is inherited by the ground-state (see Theorem 5.6), we can show that the Neumann trace of the ground-state vanishes whenever its Dirichlet trace vanishes. This allows us to extend the form domain of $H_{N}(v,w)$ and carry out the extension argument of Filonov to complete the proof of the unique continuation along the boundary in Theorem 6.1. Theorem 2.7 then follows from standard arguments.

### v2

# AMSbUmsbmn

@@ 1. Introduction
acting on the space of antisymmetric wave-functions $\wedge^{N}\mathrm{L}^{2}(I)$, where $I=(0,1)$ (or $\mathbb{R}$), and $v$ and $w$ belongs to suitable class of distributions. Such operators are used to describe interacting (spinless) electrons living in one-dimensional space and are therefore ubiquitous in quantum mechanics.
@@
In fact, our main motivation for studying the interacting operator $H_{N}(v,w)$ stems from recent advances [SPR+24, Cor25a] towards establishing a mathematically rigorous formulation of Kohn-Sham Density Functional Theory (DFT) [HK64, KS65], a highly successful and widely used approach to study the electronic structure of materials. More precisely, the class of distributional potentials studied here appeared in recent developments [SPR+24] concerning the $v$-representability problem, which, roughly speaking, consists in characterizing the set of all possible single-particle ground state densities of operators of the form $H_{N}(v,w)$ for a given class of potentials $v$. While a detailed discussion on the foundation of DFT and the $v$-representability problem goes beyond the scope of the current paper, we emphasize that the results obtained here are used to provide a completely rigorous formulation of Kohn-Sham DFT for electrons in one-dimension in a companion article [Cor25b].
@@ 2.1. Notation
We denote by $\mathrm{H}^{1}(I)$ the Sobolev space of functions $f\in\mathrm{L}^{2}(I)$ with weak derivative $\partial_{x}f\in\mathrm{L}^{2}(I)$. More generally, for $1\leq p\leq\infty$ and open $\Omega\subset\mathbb{R}^{d}$, we denote by $\mathrm{W}^{1,p}(\Omega)$ the usual Sobolev spaces of functions in $\mathrm{L}^{p}(\Omega)$ with weak gradient in $\mathrm{L}^{p}(\Omega)$, and by $\mathrm{W}^{-1,q}(\Omega)$, where $1/q+1/p=1$, the dual space of $\mathrm{W}^{1,p}(\Omega)$. In addition, we denote by $\mathrm{H}^{1/2}(\partial\Omega)$ the standard $1/2$-Sobolev (or Besov) space along the boundary $\partial\Omega$.
For a given closed subspace $B\subset\mathrm{H}^{1/2}(\partial\Omega)$, we denote by $H^{1}_{B}(\Omega)$ the Sobolev spaces of functions with boundary trace on $B$, i.e.,
$\mathrm{H}^{1}_{B}(\Omega)\coloneqq\{f\in\mathrm{H}^{1}(\Omega):\gamma f\in B\},$
where $\gamma:\mathrm{H}^{1}(\Omega)\rightarrow\mathrm{H}^{1/2}(\partial\Omega)$ is the Dirichlet boundary trace operator. The letter $B$ here stands for boundary conditions.
@@
$\mathcal{V}\coloneqq\mathrm{H}^{-1}(I;\mathbb{R})=\{v\in\mathrm{H}^{-1}(I):v(\varphi)\in\mathbb{R}\quad\mbox{for any real-valued $\varphi\in\mathrm{H}^{1}(I)$}\}.$
@@
$\mathcal{W}$ $\coloneqq\cup_{q>2}\mathrm{W}^{-1,q}(I_{2};\mathbb{R})$
$=\{w\in\mathrm{W}^{-1,q}(I_{2}):\mbox{for some }q>2\quad\mbox{and}\quad w(\varphi)\in\mathbb{R}\quad\mbox{for any real-valued $\varphi\in\mathrm{W}^{1,p}(I_{2})$}\}.$
@@
For $v\in\mathcal{V}$, $w\in\mathcal{W}$, and $B\subset\mathrm{H}^{1/2}(\partial I_{N})$ as above, we denote by $H_{N}^{B}(v,w)$ the $N$-particle Hamiltonian
$H_{N}^{B}(v,w)=-\Delta+\sum_{i\neq j}^{N}w(x_{i},x_{j})+\sum_{j=1}^{N}v(x_{i})\quad\mbox{acting on $\mathcal{H}_{N}$ with boundary conditions in $B$.}$
More precisely, $H_{N}^{B}(v,w)$ is the unique self-adjoint operator associated to the sesquilinear form
$\mathfrak{a}_{v,w}(\Psi,\Phi)\coloneqq\int_{I_{N}}\overline{\nabla\Psi(x_{1},...,x_{N})}\boldsymbol{\cdot}\nabla\Phi(x_{1},...,x_{N})\mathrm{d}x_{1}...\mathrm{d}x_{N}+v\left(\rho_{\Psi,\Phi}\right)+w(\rho^{(2)}_{\Psi,\Phi})$ (2.1)
@@
$\mathcal{Q}_{N}(B)\coloneqq\mathcal{H}_{N}\cap\mathrm{H}^{1}_{B}(I_{N}),$
@@
and $\rho^{(2)}_{\Psi,\Phi}$ is the overlapping pair density
$\rho^{(2)}_{\Psi,\Phi}(x,y)\coloneqq N(N-1)\int_{I_{N-2}}\overline{\Psi(x,y,x_{3},...,x_{N})}\Phi(x,y,x_{3},...,x_{N})\mathrm{d}x_{3}...\mathrm{d}x_{N}.$ (2.3)
@@ 2.2. Main results
As a first result, we show that the ground-state of self-adjoint realizations of $H_{N}(v,w)$ with local boundary conditions is non-degenerate. Precisely, let $\Gamma\subset\partial I_{N}$ and define
@@
Let $v\in\mathcal{V}$, $w\in\mathcal{W}$, $N\in\mathbb{N}$ and $\Gamma\subset\partial\Omega$. Then the operator $H_{N}(v,w;\Gamma)$ defined as the self-adjoint realization of (2.1) with form domain
$\mathcal{Q}_{N}(\Gamma)\coloneqq\mathrm{H}^{1}(I_{N};\Gamma)\cap\mathcal{H}_{N}$
has a unique ground-state $\Psi$ and $\Psi\neq 0$ almost everywhere.
@@
$\mathrm{H}^{1}_{\alpha}(I_{N})\coloneqq\left\{\Psi\in\mathrm{H}^{1}(I_{N}):(\gamma\Psi)\rvert_{I_{k-1}\times\{0\}\times I_{N-k}}-\alpha(\gamma\Psi)\rvert_{I_{k-1}\times\{1\}\times I_{N-k}}=0\quad\mbox{for any $1\leq k\leq N$}\right\}.$ (2.4)
@@
Then the self-adjoint realization $H_{N}^{\alpha}(v,w)$ of (2.1) with form domain
$\mathcal{Q}_{N}^{\alpha}\coloneqq\mathrm{H}^{1}_{\alpha}(I_{N})\cap\mathcal{H}_{N}$
has a unique ground state $\Psi$ and $\Psi\neq 0$ almost everywhere.
@@
Theorem 2.3 shows that periodic systems are non-degenerate for an odd number of particles, while anti-periodic systems are non-degenerate for an even number of particles. This result is optimal in the sense that there are explicit examples where the ground-state is non-degenerate if and only if condition (2.5) holds, see Appendix B. On the other hand, one can construct examples of periodic/anti-periodic systems where the ground-state is non-degenerate for any given number of particles $N$.
@@
    (Many-body interactions) We can allow a larger class of distributional interaction potentials $w$. In fact, Theorem 2.1 holds for $M$-body interactions $w_{M}$ with arbitrary $M\in\mathbb{N}$, provided that $w_{M}$ lies in suitable dual Sobolev spaces (see Section A.3). Notably, the case of two-body 3D Coulomb interactions, which could be seen as a ”six-body” interaction of the form
@@
    (More general boundary conditions) We can further extend Theorem 2.1 to a larger class of self-adjoint realizations of $H_{N}(v,w)$. Roughly speaking, the maximal class of spaces for which our proof works is given by all $\mathrm{H}^{1}_{B}(I_{N})$ where the restricted space
    $\{\Psi\rvert_{S_{N}}:\Psi\in\mathrm{H}^{1}_{B}(I_{N})\cap\mathcal{H}_{N}\}\quad\mbox{where}\quad S_{N}=\{(x_{1},...,x_{N})\in I_{N}:x_{1}<x_{2}...<x_{N}\}$
@@
Interestingly, Theorems 2.1 and 2.3 can be applied to obtain information on the whole spectrum of single-particle operators. More precisely, let $v\in\mathcal{V}$, $B\subset\mathbb{C}^{2}$ be a (closed) subspace, and denote by $h_{B}(v)=-\Delta_{B}+v$ the self-adjoint operator associated to the form
$\mathfrak{a}_{v}(\psi,\psi)=\int_{I}|\nabla\psi(x)|^{2}\mathrm{d}x+v(|\psi|^{2})\quad\mbox{with form domain $\mathcal{Q}_{1}(B)=\{\psi\in\mathrm{H}^{1}(I):(\psi(0),\psi(1))\in B\}$.}$ (2.7)
Note that $h_{B}(v)=H_{1}^{B}(v,0)$. Then by applying Theorems 2.1 and 2.3 to self-adjoint realizations of the non-interacting operator $H_{N}(v,0)$, we obtain the following result.
@@
Let $v\in\mathcal{V}$ and $h_{\alpha}(v)\coloneqq h_{B_{\alpha}}(v)$ where $B_{\alpha}=\{\beta=(\beta_{0},\beta_{1})\in\mathbb{C}^{2}:\beta_{0}-\alpha\beta_{1}=0\}$ for some $\alpha\in\mathbb{R}\setminus\{0\}$. Then $h_{\alpha}(v)$ has discrete spectrum, all eigenfunctions are almost everywhere non-vanishing, and the eigenvalues $\{\lambda_{k}\}_{k\in\mathbb{N}}$ ordered in non-decreasing order and counting multiplicity satisfy
@@
Moreover, in the case where $B$ is one of the following sets
@@
The result from Theorem 2.6 in the case of local boundary conditions (2.9) is, at least for regular potentials $v$, well-known in the theory of Sturm-Liouville operators, see, e.g., [Zet21, Theorem 1.9.2]. Nevertheless, the approach presented here seems to be completely different from previous proofs and has the important advantage that it extends to the interacting case.
@@
Moreover, if $\alpha\in\mathbb{R}\setminus\{0\}$ satisfies (2.5) and $\lambda_{1}(\Gamma)$ and $\lambda_{1}(\Gamma^{\prime})$ denote respectively the lowest eigenvalues of the self-adjoint realizations with form domain $\mathcal{Q}_{N}^{\alpha}(\Gamma)\coloneqq\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)\cap\mathcal{H}_{N}$ respectively $\mathcal{Q}_{N}^{\alpha}(\Gamma^{\prime})=\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma^{\prime})\cap\mathcal{H}_{N}$, where $\mathrm{H}^{1}_{\alpha}(I_{N};\Gamma)$ is defined in (2.6), then the same inequality holds.
@@ 2.3. Outline of the proofs
The proof of Theorems 2.1 and 2.3 is carried out in Section 4 and relies on two main steps. The first step consists in proving a Perron-Frobenius (PF) theorem for Schrödinger operators of the form $\Delta+V$ for a large class of distributional potentials $V$ (see Theorem 4.6). In other words, we show that the ground-state of such operators is non-degenerate and strictly positive almost everywhere. This step relies on well-known semigroup techniques and a perturbative argument borrowed from Reed and Simon [RS78][Section XII.12] (see also [GJ70, Gol77, FS75, Ges84] for similar arguments). Combining this result with a density result for dual Sobolev spaces (Lemma 3.11), we can then establish a PF theorem for Schrödinger operators of the form $H_{N}(v,w)=-\Delta+\sum_{j}v_{j}+\sum_{j\neq i}w_{i,j}$ under a suitable positivity assumption on their form domain. Unfortunately, these results only apply to the operator without Fermi statistics, i.e., without antisymmetry restrictions on the wave-function. Indeed, whenever antisymmetry is present the wave-function must have both positive and negative values and can not be almost everywhere strictly positive. In particular, as highlighted in [RS78, page 207], PF results cannot be applied to electronic systems.
@@
Remark 2.8 (Jordan-Wigner transform and Bose-Fermi duality in 1D).
As brought to my attention by an anonymous referee, it turns out that the strategy described above is related to the Jordan-Wigner (JW) transform [JW28] or bosonization approach, which is apparently well-known in the physics community. More precisely, while this transform is usually applied to establish an equivalence between spin systems and 1D lattice fermions (see [Gia04, Chapter 6]), a continuum analog has been used in a celebrated work of Girardeau [Gir60] to establish the existence of a Fermi-Bose map. This continuum JW transform (or sign map) of Girardeau can be shown to be equivalent to the simplex reduction strategy used here, see Remark 4.9. Nevertheless, we emphasize that [Gir60] mostly deal with the case of Bosons with purely hard-core interactions (the Tonks-Girardeau gas), for which the equivalent fermionic system is non-interacting.
@@
In Section 4.4 we present the proof of Theorem 2.6. This result is an immediate application of Theorems 2.1 and 2.3, and a simple lemma that connects the boundary conditions of the non-interacting operator $H_{N}(v,0)$ to the boundary conditions of the single-particle operator $h(v)=-\Delta+v$.
@@
The proof of Theorem 2.7 is inspired by ideas from previous works on similar eigenvalue inequalities [Fil05, GM09, BRS18, Roh21]. More precisely, the core idea of the proof is to apply the beautiful domain extension argument of Filonov [Fil05] to establish a unique continuation result along the boundary (cf. Theorem 5.1). Here, however, we face two fundamental problems when trying to apply this argument. First, it relies on unique continuation results which are not available in the case of distributional potentials. Second, and perhaps more critical, is the fact that the non-local boundary conditions are such that the faces of $I_{N}$ are ”glued” to each other, and therefore, no extension argument is possible.
@@
The first problem can be overcome by using the unique continuation property of the ground-state guaranteed by the PF theorem. Note that this property is only proven for the ground-state, which explains why we can only obtain inequalities for the lowest eigenvalue of $H_{N}(v,w)$. To deal with the second difficulty, we rely on a distributional definition of the Neumann trace (or normal derivative). While this definition is the same as in previous works, see, e.g., [GM09] where the Laplacian with non-local (Robin) boundary conditions is considered, we emphasize that the situation here is more delicate for two reasons. First, the non-local boundary conditions here are imposed directly in the form domain of the operator, as opposed to adding a non-local form along the boundary as done in [GM09]. Second, an eigenfunction $\Psi$ of $H_{N}(v,w)$ has, in general, no more regularity than $\mathrm{H}^{1}$. In particular, its Laplacian has no more regularity than $\mathrm{H}^{-1}$, and therefore, the usual definition of the weak Neumann trace, which requires $\Delta\Psi$ to be at least in $\mathrm{H}^{s}$ for s>1/2s>-1/2, see [GM09, (3.17)], is not meaningful.
@@
Fortunately, despite this lack of regularity, we can still show that a meaningful sense can be given to the Neumann trace of eigenfunctions along any open subset of the boundary where said eigenfunction vanish. This result crucially relies on the locality of the operator $H_{N}(v,w)$, a property that indirectly underlies many of the arguments used here. We then obtain an explicitly formula for this weak Neumann trace that could be new and of independent interest; see Lemma 5.4. By combining this formula with a positivity property of the boundary conditions, which is inherited by the ground-state (see Theorem 4.8), we can show that the Neumann trace of the ground-state vanishes whenever its Dirichlet trace vanishes. This allows us to extend the form domain of $H_{N}(v,w)$ and carry out the extension argument of Filonov to obtain a contradiction, thereby completing the proof of the boundary unique continuation result in Theorem 5.1. Theorem 2.7 then follows from standard arguments.